RC meeting and Run Plan, Tuesday February 14th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Report


  • Ran with ~30 uA beam current most of yesterday, collecting 0.7C
  • At 9 am the accelerator went down due to many RF issues. Will be down until 8 pm.
  • BBCal block now look good - https://logbooks.jlab.org/entry/4130863
  • GEM experts are in the hall working on SBS GEMs


Current Status

  • Running 45 uA on He3

Run Plan

  • Continue running 45 uA on He3
  • Accelerator will prioritize beam and not have any maintenance days this week
  • Hall B wants to change the IHWP on Friday

Data Collection and Polarization Status

  • 30.9C of uncorrected charge for this kinematic (35.8C of corrected charge)
  • Last NMR 50.0%%

Other Topics/Status Updates

  • Anything else?
  • Bill wants to change the raster settings after some analysis

Requested Accesses to the Hall

  • Jack would like to collect some cables
  • GEM team need like a 3-4hour access for the SBS GEMs DAQ troubleshooting